OpenStudy (anonymous):

Translate the following statement into an appropriate algebraic equation: The number of dogs (d) in the animal shelter is twice the sum of 4 and the number of cats (c).

OpenStudy (anonymous):

Okay, so it would be d = , that's the first part The sum of 4 and c just means 4+c. So d = 2 times 4 + c, now can you simplify that?

OpenStudy (anonymous):

2(4+c)

OpenStudy (anonymous):

There we go. So your answer would be d = 2(4 + c), or d = 8 + 2c, depending on how it needs to be written. I think the first one would be a little bit more appropriate, considering the question.

OpenStudy (anonymous):

ok thank you c: